ILPNET (Expenditure claim for durable equipment and consumables, specified in Annex 1)

The undersigned

(name in print)

representative of

(institution)

claim following expenses:

1. Communication costs:

(amount in ecu, max. 200 ecu per year)

2. Purchase of books and periodicals: (for Central/Eastern nodes only)

(amount in ecu, max. 700 ecu per year)

3. Networking costs:

(for Central/Eastern nodes only)

(amount in ecu, max. 3300 ecu in total)

I have read Annex 1 of the contract and I certify that this claim is consistent with the Annex. The documentation about these costs (except for the communication costs) can be made available to the CEC upon request.

Signed Date